OpenStudy (anonymous):

What is (-7)-(-4)= plzz explain with great details

OpenStudy (anonymous):

-7 - (-4) is = -7 +4... so when you subtract two negatives, you change one of the negatives into a positive

OpenStudy (anonymous):

-3 because a negative times a negative gives you a positive. so -7+4=-3

OpenStudy (anonymous):

Minus a negative makes it positive, as a rule (-7)-(-4) = (-7)+4 = -3

OpenStudy (anonymous):

Ohhhh. That's the part I always get confused on

OpenStudy (anonymous):

Just remember two negatives don't make a negative, it makes a positive

OpenStudy (cggurumanjunath):